OpenStudy (anonymous):

Name the property the equation illustrates. 9 + (–9) = 0 a Identity Property of Addition b Addition Property of 0 c Inverse Property of Addition d Inverse Property of Multiplication

OpenStudy (anonymous):

Commutative Property of Addition a + b = b + a ex:) 2 + 3 = 3 + 2

OpenStudy (anonymous):

Commutative Property of Multiplication a • b = b • a ex:) 2 • ( 3 ) = 3 • ( 2 )

OpenStudy (anonymous):

Associative Property of Addition a + ( b + c ) = ( a + b ) + c ex:) 2 + ( 3 + 4 ) = ( 2 + 3 ) + 4

OpenStudy (anonymous):

Associative Property of Multiplication a • ( b • c ) = ( a • b ) • c ex:) 2 • ( 3 • 4 ) = ( 2 • 3 ) • 4

OpenStudy (anonymous):

Distributive Property a • ( b + c ) = a • b + a • c ex:) 2 • ( 3 + 4 ) = 2 • 3 + 2 • 4

OpenStudy (anonymous):

Additive Identity Property a + 0 = a ex:) 3 + 0 = 3

OpenStudy (anonymous):

Multiplicative Identity Property a • 1 = a ex:) 3 • 1 = 3

OpenStudy (anonymous):

Additive Inverse Property a + ( -a ) = 0 ex:) 3 + (-3) = 0

OpenStudy (anonymous):

Zero Property a • 0 = 0 ex:) 5 • 0 = 0

OpenStudy (anonymous):

so @anaivillegas which is it? what property? i gave you an example.

OpenStudy (anonymous):

i think additive inverse property

OpenStudy (anonymous):

Additive Inverse Property yes precisely :) good job ;D

OpenStudy (anonymous):

thank you